OpenStudy (anonymous):

If f(x)=4x+2, find f(3) f(3)= ?

OpenStudy (anonymous):

You have to just plug in

OpenStudy (anonymous):

when f(#) Its the same as f(x) because x is just a variable and so if its saying f(3) you just have yo plug in for x and solve

OpenStudy (kkutie7):

all this means is that every where there is an x you put a 3 =) then solve

OpenStudy (anonymous):

x=4?

OpenStudy (anonymous):

No you plug in x and solve

OpenStudy (kkutie7):

no \[f(3)=4(3)+2\]

OpenStudy (anonymous):

\[f(x)=f(3)\] they are the same so like kkutie7 placed it yea

OpenStudy (anonymous):

you just use the number they give you to plug in and then continue to solve the equation using PEMDAS

OpenStudy (anonymous):

i dont know how to solve it tho

OpenStudy (anonymous):

@BinkyBoo

OpenStudy (anonymous):

By using PEMDAS when you have f(x) its like having just another variable. Variables are just place holders for numbers that are not known yet, which if usually what you need to find

OpenStudy (anonymous):

f=2 @BinkyBoo ?

OpenStudy (anonymous):

Its like this for example Find f(-15) for \[f(x)=9-4x\] First you plug in \[f(-5)=9-4(-5)\] Then You Start PEMDAS Since there is parenthesis but only one variable in them you move to exponents no exponents move to multiplication \[f(-5)=9-(-20)\] Due to 4(-5)= -20 Then Two negatives multiplying together make a positive making the equation \[f(-5)=9+20\] Then Addition \[f(-5)=29\] And thats it!

OpenStudy (anonymous):

whoops in the beginning that 15 was suppose to be a five I'm so sorry

OpenStudy (anonymous):

is it 44 @BinkyBoo ?

OpenStudy (anonymous):

Walk me through your steps

OpenStudy (anonymous):

everything you did except 15

OpenStudy (anonymous):

well i mixed up the problem a bit like you don't have the negative signs and everything

OpenStudy (anonymous):

but its correct?

OpenStudy (anonymous):

no not for your question